# Characterizing the Burst Error Correction Ability of Quantum Cyclic Codes

### v1

@@ IV The True Burst Error Correction Ability of Quantum Reed-Solomon Codes
Similar to classical RS codes [7], quantum RS codes are efficient to correct quantum burst errors [51]. Moreover, there exists a lower bound for the burst error correction ability of quantum RS codes [15, 7]. However, this lower bound is not tight and does not give the true burst error correction ability of quantum RS codes. In this section, we give a polynomial-time algorithm to determine that.
Let ${C}_{RS}=[n=q^{m}-1,k]_{q^{m}}$ be a classical narrow sense RS code such that $n\leq 2k$. Then we know that ${C}_{RS}^{\bot}\subseteq{C}_{RS}$ and we can construct a ${Q}_{RS}=[[n,2k-n]]_{q^{m}}$ quantum RS code by using the CSS construction [52]. Let $[{C}_{RS}]=[mn,mk]_{q}$ be the $q$-ary image of the RS code ${C}_{RS}$. We use the self-dual basis of $\mathbb{F}_{q^{m}}$ over $\mathbb{F}_{q}$ so that the dual-containing relationship can be maintained in the $q$-ary extension of ${C}_{RS}$ [53, 54].
@@
[55] Let $\{\alpha_{1},\cdots,\alpha_{m}\}$ be a self-dual basis of $\mathbb{F}_{q^{m}}$ over $\mathbb{F}_{q}$, i.e., $\textrm{Tr}(\alpha_{i}\alpha_{j})=\delta_{ij}$ for all $1\leq i,j\leq m$. Let Let $[{C}_{RS}]$ and $[{C}_{RS}]^{\bot}$ be the $q$-ary images of ${C}_{RS}$ and ${C}_{RS}^{\bot}$ under the basis $\{\alpha_{1},\cdots,\alpha_{m}\}$, respectively. If ${C}_{RS}^{\bot}\subseteq{C}_{RS}$, then $[{C}_{RS}]^{\bot}\subseteq[{C}_{RS}]$ and $[{C}_{RS}]^{\bot}$ is the dual of $[{C}_{RS}]$.
@@
Let $C=[n,k]_{q^{m}}$ be a $q^{m}$-ary RS code. Denote $\mathbf{D}=[C]$ by the $q$-ary image of $C$ under a self-dual basis $\{\alpha_{1},\cdots,\alpha_{m}\}$. Let $\mu_{1}$ and $\mu_{2}$ be any two codewords of $C$. Let $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$ be any two codewords of $\mathbf{D}$. Map $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$ to two vectors $\nu_{1},\nu_{2}\in\mathbb{F}_{q^{m}}^{n}$ by using the basis $\{\alpha_{1},\cdots,\alpha_{m}\}$. If $\mu_{1}\neq\mu_{2}$ and $\mathbf{v}_{1}\neq\mathbf{v}_{2}$, then $[\mu_{1}]\neq[\mu_{2}]$ and $\nu_{1}\neq\nu_{2}$.
@@
By using Lemma 6 and the CSS construction, we can construct a $[{Q}_{RS}]=[[mn,2mk-mn]]_{q}$ quantum code and we call it the image of the quantum RS code ${Q}_{RS}$. Denote by $\hbar=\lfloor(n-k)/2\rfloor$. Then the $q$-ary image code $[{Q}_{RS}]$ can correct any quantum burst error of length $(\hbar-1)m+1$ or less [7, 15].
@@
However, the burst error correction ability of $[{Q}_{RS}]$ is a lower bound and does not give the true burst error correction ability of $[{Q}_{RS}]$. In this work, we give a polynomial-time algorithm to calculate the true burst error correction ability of quantum RS codes.
Let $B_{\imath}=[\imath,\imath+\hbar]$, where $11\leq\imath\leq n-2\hbar-$ 1. Let $\mathcal{M}^{(\hbar+1)}$ be the $(r-\hbar-1)\times(n-\hbar-1)$ matrix formed by deleting the last $\hbar+1$ rows and last $\hbar+1$ columns of the parity-check matrix $\mathbf{H}$ of the RS code. Let $\mathcal{M}^{(\hbar+1)}_{B_{\imath}}$ be the subblock of $\hbar+1$ consecutive columns of $\mathcal{M}^{(\hbar+1)}$ indexed by $B_{\imath}$. Denote the set of columns of $\mathcal{M}^{(\hbar+1)}_{B_{\imath}}$ by $\{\alpha_{B_{\imath}}^{(1)},\cdots,\alpha_{B_{\imath}}^{(\hbar+1)}\}$. We have the following result about $\mathcal{M}^{(\hbar+1)}_{B_{\imath}}$.
@@
The rank of each $\mathcal{M}^{(\hbar+1)}_{B_{\imath}}$ satisfies $\hbar-1\leq rank(\mathcal{M}^{(\hbar+1)}_{B_{\imath}})\leq\hbar$, where $11\leq\imath\leq n-2\hbar-$ 1.
@@
^$\widehat{\boxtimes}_{\mathcal{M}^{(\hbar+1)}_{B_{\imath}}}=\{(e,f)|(e,f)\in\boxtimes_{\mathcal{M}^{(\hbar+1)}_{B_{\imath}}},\mathbf{G}e^{T}\neq\mathbf{G}f^{T}\}.$ (10)
@@
Let ${C}=[n=q^{m}-1,k]_{q^{m}}$ be a classical RS code satisfying the dual containing relationship, i.e., ${C}^{\bot}\subseteq{C}$. Let $\hbar=\lfloor(n-k)/2\rfloor$ and let $B_{\imath}=[\imath,\imath+\hbar]$, where $11\leq\imath\leq n-2\hbar-$ 1. Let $\mathcal{L}$ be largest integer $\ell$ such that there does not exist $(e,f)\in\widehat{\boxtimes}_{\mathcal{M}^{(\hbar+1)}_{B_{\imath}}}$ with
@@
There exists a $[[n,2k-n]]_{q^{m}}$ quantum RS code $Q$ which can correct any quantum burst error of length $\mathcal{L}$ or less.
@@
4:     ///// / Fail to construct a quantum RS code.
@@
10:     ///// /Do Gaussian elimination to $\mathcal{M}^{(\hbar+1)}_{B_{\imath}}$.
@@
16:         ^$\widehat{\boxtimes}_{\mathcal{M}^{(\hbar+1)}_{B_{\imath}}}=\{(e,f)|(e,f)\in\boxtimes_{\mathcal{M}^{(\hbar+1)}_{B_{\imath}}},$
@@
TABLE III: The True Burst Error Correction Ability of Quantum RS Codes.
[…]
$4$ $15$ $5$ $8$ $5$ $10$
$1$ $12$ $9$ $14$
$5$ $31$ $23$ $7$ $6$ $10$
$21$ $8$ $6$ $12$
@@ V Error-Trapping Decoder of Quantum Cyclic Codes
In this section, we present the quantum error-trapping decoder (QETD) for QCCs, and we use the CSS-type QCCs to illustrate the proceess of decoding. Let ${C}=[n,k]$ be a classical cyclic code over $\mathbb{F}_{2}$ such that ${C}^{\bot}\subseteq{C}$. Denote the generator polynomial of ${C}$ by $\mathbf{g}(x)$, and denote the parity-check matrix of ${C}$ by $\mathbf{H}$. Let $\mathcal{Q}=[[n,\mathcal{K}=2k-n]]$ be a QCC constructed from ${C}$ by using the CSS construction. Suppose that $\mathcal{Q}$ can correct any quantum burst error of length $\mathcal{L}$ or less. According to the quantum Reiger bound in Theorem 1, we have $\mathcal{L}\leq(n-\mathcal{K})/4$. In this section we present the quantum error-trapping decoder for QCCs to decode any quantum burst error of length up to $\mathcal{L}$. Similar to the classical error-trapping decoder, we show that QETD can also correct additional quantum burst errors of length $\mathcal{L}<l\leq(n-\mathcal{K})/2$. Such burst errors belong to the coset leaders of ${C}$. In addition, we will show that QETD can also decode degenerate burst errors belong to the coset of ${C}^{\bot}$.
@@
and denote by $\mathbf{S}(x)=s_{0}+s_{1}x+\cdots+s_{r-1}x^{r-1}$, where $r=n-k$. Moreover, the syndrome polynomial $\mathbf{S}(x)$ is equal to the remainder of dividing $\mathbf{e}(x)$ by the generator $\mathbf{g}(x)$, i.e.,
@@
Let $\mathbf{e}=(e_{0},e_{1},\cdots,e_{n-1})$ be a correctable burst error of length $2\leq\ell\leq(n-\mathcal{K})/2=n-k$. It is natural to suppose that the burst length of $\mathbf{e}$ is less than or equal to $n-k$. If the quantum burst error $\mathbf{e}$ is confined to the $n-k$ low-order positions, then $\mathbf{e}(x)=e_{0}+e_{1}x+\cdots+e_{n-k-1}x^{n-k-1}$. According to Eq. (13), we know that $\mathbf{e}(x)=u(x)\mathbf{g}(x)+\mathbf{S}(x)$. Since the degree of $\mathbf{e}(x)$ is less than $n-k$, we have $\mathbf{S}(x)=\mathbf{e}(x)=e_{0}+e_{1}x+\cdots+e_{n-k-1}x^{n-k-1}$. That is to say we can derive the burst error $\mathbf{e}(x)$ from the syndrome directly as long as $\mathbf{e}$ is confined to the $n-k$ low-order positions. Moreover, all the burst errors that confined to the $n-k$ low-order positions range over all the possible $q^{n-k}$ error syndromes for ${C}$. This possibility is the key for the decoding of QCCs.
@@
If the burst error $\mathbf{e}$ is not confined to the $n-k$ low-order positions, we cyclically shift the syndrome $\mathbf{S}$ by a certain number of times to capture $\mathbf{e}$. If $0\leq\textbf{bl}(\mathbf{e})\leq\ell$, then $\mathbf{e}$ or an element in the coset $\mathbf{e}+{C}^{\bot}$ can be captured in the $n-k$ low-order positions. If $\ell+1\leq\textbf{bl}(\mathbf{e})\leq n-k$, then $\mathbf{e}$ may exceed the burst error correction ability of $\mathcal{Q}$. However, we can still correct such burst error if it is a coset leader of $C$ or it belongs to the coset of ${C}^{\bot}$. Before doing so, we first cyclically shift the syndrome $n$ times to determine the shortest burst that is confined to the $n-k$ low-order positions as in [7, 56]. Then we give a circuit for QETD in Fig. 1 and present the whole process of QETD as follows.
@@
    Shift the syndrome register with Gate 2 on and a clock starts to count simultaneously. We use a counter $s$ to record the length $t$ of continuous zeros in the leftmost stages of the syndrome register, and $s$ is set to 00 before the clock. If the length $t$ of continuous zeros in the leftmost stages of the syndrome register is larger than the previous one, $s$ is updated by the length $t$ in the current clock. After $n$ clocks, the counter $s$ is the longest length of continuous zeros in the leftmost stages of the syndrome register. Then $z\equiv r-s$ is the shortest length of burst errors that appear in the $z$ rightmost stages of the syndrome register.
@@
    Shift the syndrome register with Gate 2 on. As soon as the $r-z$ leftmost stages of the syndrome register contain all zeros after the $i$th shift for $10\leq i\leq n-$ 1, the burst error is confined in the $z$ rightmost stages. Then Gate 2 is turned off.
@@
    The syndrome register is shifted by $z$ times with Gate 3 on. Then the error burst is confined to the error sequence register and Gate 3 is turned off. We need to continue to shift the error register so that the error burst is put in the right positions. With Gate 3 on, the error sequence register is cyclically shifted by $[(n-z-i)\mod n]$ times.
The decoding circuit in Fig. 1 is easy to be implemented by using the linear shift register and the time complexity is linear. On the other hand, for the purpose of numerical simulations, we give a simplified QETD algorithm for QCCs in Algorithm 3.
@@
3:Initialization: $\textbf{Z}=+\infty$, ${v}=0$;
4:///// / Calculate the shortest burst that is confined in the right most stages of the syndrome.
@@
12:///// / Shift the error sequence to the right position.
13:$\textbf{e}(x)=x^{-v}(x^{v}\textbf{S}(x)\mod\textbf{g}(x))\mod(x^{n}-1);$
@@
The QETD in Algorithm 3 can decode all correctable burst errors of length $\ell\leq(n-\mathcal{K})/2$ and is quantum maximum likelihood decoding according to [47]. Furthermore, the QETD can also decode degenerate errors belong to the coset of $C^{\bot}$. In addition, QETD in Algorithm 3 is also available for Hermitian-ype QCCs. In Table IV, we list the numerical results of several QCCs for correcting quantum burst errors. Denote $\mathcal{N}_{D}$, $\mathcal{N}_{0}$, and $\mathcal{N}$ by the numbers of all correctable burst errors, all correctable nondegenerate burst errors and total errors of length $n\leq(n-\mathcal{K})/2$, respectively. We exhaustively traverse all the burst errors of length $\ell\leq(n-\mathcal{K})/2$ and count the numbers of burst errors that are successfully decoded by QETD. It is shown in Table IV that QETD can correct more degenerate burst errors than nondegenerate ones. As the code length grows, the ratio $\mathcal{N}_{D}/\mathcal{N}_{0}$ becomes extremely larger and thus QETD can correct much more degenerate burst errors than nondegenerate ones. As an example, nondegenerate burst errors that can be decoded by the $Q=[[29,1]]$ QCC are only $4.4\%$ of the total burst errors of length $\ell\leq 12$. However, $Q$ can decode degenerate burst errors up to $26.9\%$ of the total burst errors of length $\ell\leq 12$.
@@
$[[5,1]]$ $15$ $15$ $51$ $29.4\%$ $29.4\%$ $1$ $g=(\textbf{1}^{2}\textbf{2}^{1}\textbf{1}^{0})$
$[[7,1]]$ $72$ $57$ $255$ $28.2\%$ $22.3\%$ $1.26$ $g=(\textbf{1}^{3}\textbf{1}^{1}\textbf{1}^{0})$
@@
[…]

### v2

@@ IV The True Burst Error Correction Ability of Quantum Reed-Solomon Codes
As with classical RS codes [37], quantum RS codes are effective at correcting both quantum random errors [42, 48] and quantum burst errors [24]. Moreover, there exists a lower bound for the burst error correction ability of quantum RS codes [45, 37]. However, this lower bound is not tight and does not give the true burst error correction ability of quantum RS codes. In this section, we give a polynomial-time algorithm to determine that.
Let ${C}_{RS}=[n=q^{m}-1,k]_{q^{m}}$ be a classical narrow sense RS code such that $n\leq 2k$. Then we know that ${C}_{RS}^{\bot}\subseteq{C}_{RS}$ and we can construct a ${Q}_{RS}=[[n,2k-n]]_{q^{m}}$ quantum RS code by using the CSS construction [46]. Let ${C}_{RS}]=[mn,mk]_{q}$ be the $q$-ary image of the RS code ${C}_{RS}$. We use the self-dual basis of $\mathbb{F}_{q^{m}}$ over $\mathbb{F}_{q}$ so that the dual-containing relationship can be maintained in the $q$-ary extension of ${C}_{RS}$ [38, 9].
@@
[1] Let $\{\alpha_{1},\cdots,\alpha_{m}\}$ be a self-dual basis of $\mathbb{F}_{q^{m}}$ over $\mathbb{F}_{q}$, i.e., $\textrm{Tr}(\alpha_{i}\alpha_{j})=\delta_{ij}$ for all $1\leq i,j\leq m$. Let ${C}_{RS}]$ and ${C}_{RS}]^{\bot}$ be the $q$-ary images of ${C}_{RS}$ and ${C}_{RS}^{\bot}$ under the basis $\{\alpha_{1},\cdots,\alpha_{m}\}$, respectively. If ${C}_{RS}^{\bot}\subseteq{C}_{RS}$, then ${C}_{RS}]^{\bot}\subseteq[{C}_{RS}]$ and ${C}_{RS}]^{\bot}$ is the dual of ${C}_{RS}]$.
@@
Let $C=[n,k]_{q^{m}}$ be a $q^{m}$-ary RS code. Denote $\mathbf{D}=[C]$ by the $q$-ary image of $C$ under a self-dual basis $\{\alpha_{1},\cdots,\alpha_{m}\}$. Let $\mu_{1}$ and $\mu_{2}$ be any two codewords of $C$. Let $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$ be any two codewords of $\mathbf{D}$. Map $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$ to two vectors $\nu_{1},\nu_{2}\in\mathbb{F}_{q^{m}}^{n}$ by using the basis $\{\alpha_{1},\cdots,\alpha_{m}\}$. If $\mu_{1}\neq\mu_{2}$ and $\mathbf{v}_{1}\neq\mathbf{v}_{2}$, then $\mu_{1}]\neq[\mu_{2}]$ and $\nu_{1}\neq\nu_{2}$.
@@
By using Lemma 6 and the CSS construction, we can construct a ${Q}_{RS}]=[[mn,2mk-mn]]_{q}$ quantum code and we call it the image of the quantum RS code ${Q}_{RS}$. Denote by $\hbar=\lfloor(n-k)/2\rfloor$. Then the $q$-ary image code ${Q}_{RS}]$ can correct any quantum burst error of length $(\hbar-1)m+1$ or less [37, 45].
@@
However, the burst error correction ability of ${Q}_{RS}]$ is a lower bound which does not give the true burst error correction ability of ${Q}_{RS}]$. In this work, we give a polynomial-time algorithm to calculate the true burst error correction ability of quantum RS codes.
Let $B_{\imath}=[\imath,\imath+\hbar-1]$, where $1\leq\imath\leq n-2\hbar-1$. Let $\mathcal{M}^{(\hbar+1)}$ be the $(r-\hbar-1)\times(n-\hbar-1)$ matrix formed by deleting the last $\hbar+1$ rows and last $\hbar+1$ columns of the parity-check matrix $\mathbf{H}$ of the RS code. Let $\mathcal{M}^{(\hbar+1)}_{B_{\imath}}$ be the subblock of $\hbar+1$ consecutive columns of $\mathcal{M}^{(\hbar+1)}$ indexed by $B_{\imath}$. Denote the set of columns of $\mathcal{M}^{(\hbar+1)}_{B_{\imath}}$ by $\{\alpha_{B_{\imath}}^{(1)},\cdots,\alpha_{B_{\imath}}^{(\hbar+1)}\}$. We have the following result about $\mathcal{M}^{(\hbar+1)}_{B_{\imath}}$.
@@
The rank of each $\mathcal{M}^{(\hbar+1)}_{B_{\imath}}$ satisfies $\hbar-1\leq rank(\mathcal{M}^{(\hbar+1)}_{B_{\imath}})\leq\hbar$, where $1\leq\imath\leq n-2\hbar-1$.
@@
$\widehat{\boxtimes}_{\mathcal{M}^{(\hbar+1)}_{B_{\imath}}}=\{(e,f)|(e,f)\in\boxtimes_{\mathcal{M}^{(\hbar+1)}_{B_{\imath}}},\mathbf{G}e^{T}\neq\mathbf{G}f^{T}\}.$ (10)
@@
Let ${C}=[n=q^{m}-1,k]_{q^{m}}$ be a classical RS code satisfying the dual containing relationship, i.e., ${C}^{\bot}\subseteq{C}$. Let $\hbar=\lfloor(n-k)/2\rfloor$ and let $B_{\imath}=[\imath,\imath+\hbar]$, where $1\leq\imath\leq n-2\hbar-1$. Let $\mathcal{L}$ be largest integer $\ell$ such that there does not exist $(e,f)\in\widehat{\boxtimes}_{\mathcal{M}^{(\hbar+1)}_{B_{\imath}}}$ with
@@
There exists a $n,2k-n]]_{q^{m}}$ quantum RS code $Q$ which can correct any quantum burst error of length $\mathcal{L}$ or less.
@@
4:  //// Fail to construct a quantum RS code.
@@
10:  ////Do Gaussian elimination to $\mathcal{M}^{(\hbar+1)}_{B_{\imath}}$.
@@
16:   $\widehat{\boxtimes}_{\mathcal{M}^{(\hbar+1)}_{B_{\imath}}}=\{(e,f)|(e,f)\in\boxtimes_{\mathcal{M}^{(\hbar+1)}_{B_{\imath}}}$,
@@
TABLE III: The true burst error correction ability of quantum RS codes. The powers of the self-dual basis of $\mathbb{F}_{2^{m}}$ over $\mathbb{F}_{2}$ is denoted by $\mathrm{SDB}_{m}$.
$m$ $n$ $\mathrm{SDB}_{m}$ $k$ $\mathcal{L}$
$\ell_{L}$ in
[45, 37]
QRB in
Theorem 1
44 $15$ {6,9,11,146,9,11,14} 55 88 55 $10$
11 $12$ 99 $14$
55 $31$ {3,5,11,22,243,5,11,22,24} $23$ 77 66 $10$
2121 99 66 $12$
1919 $12$ $11$ $15$
@@ V Error-Trapping Decoder of Quantum Cyclic Codes
In this section, we present the quantum error-trapping decoder (QETD) for QCCs, and we use the CSS-type QCCs to illustrate the proceess of decoding. Let C=[n,k]{C}=[n,k] be a classical cyclic code over $\mathbb{F}_{2}$ such that ${C}^{\bot}\subseteq{C}$. Denote the generator polynomial of C{C} by $\mathbf{g}(x)$, and denote the parity-check matrix of C{C} by $\mathbf{H}$. Let $\mathcal{Q}=[[n,\mathcal{K}=2k-n]]$ be a QCC constructed from C{C} by using the CSS construction. Suppose that $\mathcal{Q}$ can correct any quantum burst error of length $\mathcal{L}$ or less. According to the quantum Reiger bound in Theorem 1, we have $\mathcal{L}\leq(n-\mathcal{K})/4$. In this section we present the quantum error-trapping decoder for QCCs to decode any quantum burst error of length up to $\mathcal{L}$. Similar to the classical error-trapping decoder, we show that QETD can also correct additional quantum burst errors of length $\mathcal{L}<l\leq(n-\mathcal{K})/2$. Such burst errors belong to the coset leaders of C{C}. In addition, we will show that QETD can also decode degenerate burst errors that belong to the coset of ${C}^{\bot}$.
@@
and denote by $\mathbf{S}(x)=s_{0}+s_{1}x+\cdots+s_{r-1}x^{r-1}$, where r=nkr=n-k. Moreover, the syndrome polynomial $\mathbf{S}(x)$ is equal to the remainder of dividing $\mathbf{e}(x)$ by the generator $\mathbf{g}(x)$, i.e.,
@@
Let $\mathbf{e}=(e_{0},e_{1},\cdots,e_{n-1})$ be a correctable burst error of length $2\leq\ell\leq(n-\mathcal{K})/2=n-k$. It is natural to suppose that the burst length of $\mathbf{e}$ is less than or equal to nkn-k. If the quantum burst error $\mathbf{e}$ is confined to the nkn-k low-order positions, then $\mathbf{e}(x)=e_{0}+e_{1}x+\cdots+e_{n-k-1}x^{n-k-1}$. According to Eq. (13), we know that $\mathbf{e}(x)=u(x)\mathbf{g}(x)+\mathbf{S}(x)$. Since the degree of $\mathbf{e}(x)$ is less than nkn-k, we have $\mathbf{S}(x)=\mathbf{e}(x)=e_{0}+e_{1}x+\cdots+e_{n-k-1}x^{n-k-1}$. That is to say we can derive the burst error $\mathbf{e}(x)$ from the syndrome directly as long as $\mathbf{e}$ is confined to the nkn-k low-order positions. Moreover, all the burst errors that confined to the nkn-k low-order positions range over all the possible $q^{n-k}$ error syndromes for C{C}. This possibility is the key for the decoding of QCCs.
@@
If the burst error $\mathbf{e}$ is not confined to the nkn-k low-order positions, we cyclically shift the syndrome $\mathbf{S}$ by a certain number of times to capture $\mathbf{e}$. If $0\leq\textbf{bl}(\mathbf{e})\leq\ell$, then $\mathbf{e}$ or an element in the coset $\mathbf{e}+{C}^{\bot}$ can be captured in the nkn-k low-order positions. If $\ell+1\leq\textbf{bl}(\mathbf{e})\leq n-k$, then $\mathbf{e}$ may exceed the burst error correction ability of $\mathcal{Q}$. However, we can still correct such burst error if it is a coset leader of $C$ or it belongs to the coset of ${C}^{\bot}$. Before doing so, we first cyclically shift the syndrome $n$ times to determine the shortest burst that is confined to the nkn-k low-order positions as in [37, 22]. Then we give a circuit for QETD in Fig. 1 and present the whole process of QETD as follows.
@@
    Shift the syndrome register with Gate 2 on and a clock starts to count simultaneously. We use a counter $s$ to record the length $t$ of continuous zeros in the leftmost stages of the syndrome register, and $s$ is set to 0 before the clock. If the length $t$ of continuous zeros in the leftmost stages of the syndrome register is larger than the previous one, $s$ is updated by the length $t$ in the current clock. After $n$ clocks, the counter $s$ is the longest length of continuous zeros in the leftmost stages of the syndrome register. Then $z\equiv r-s$ is the shortest length of burst errors that appear in the $z$ rightmost stages of the syndrome register.
@@
    Shift the syndrome register with Gate 2 on. As soon as the rzr-z leftmost stages of the syndrome register contain all zeros after the $i$th shift for $0\leq i\leq n-1$, the burst error is confined in the $z$ rightmost stages. Then Gate 2 is turned off.
@@
    The syndrome register is shifted by $z$ times with Gate 3 on. Then the error burst is confined to the error sequence register and Gate 3 is turned off. We need to continue to shift the error register so that the error burst is put in the right positions. With Gate 3 on, the error sequence register is cyclically shifted by $(n-z-i)\mod n]$ times.
The decoding circuit in Fig. 1 is easy to be implemented by using the linear shift register. In steps (2) and (3), the shift registers both run in $O(n)$ time. In step (4), the syndrome register also runs in $O(n)$ time. Therefore the total time complexity of the QETD algorithm is linear. On the other hand, for the purpose of numerical simulations, we give a simplified QETD algorithm for QCCs in Algorithm 3. It should be noted that the time complexity of Algorithm 3 is indeed $O(n^{2})$ which is higher than the complexity of the circuit level decoder in Fig. 1. But Algorithm 3 is easier to be simulated in computers by using high level programming languages such as C/C++, Magma and Python, etc.
@@
3:Initialization: $\textbf{Z}=+\infty$, v=0{v}=0;
4://// Calculate the shortest burst that is confined in the right most stages of the syndrome.
@@
12://// Shift the error sequence to the right position.
[…]
@@
The QETD in Algorithm 3 can decode all correctable burst errors of length $\ell\leq\mathcal{L}$ which indicates the burst error correction ability of the QCCs. Moreover, the QETD algorithm can also decode all correctable burst errors of length up to $(n-\mathcal{K})/2$ that are the coset leaders, which have the shortest burst length in the coset. Therefore, the QETD algorithm is quantum maximum likelihood decoding of nondegenerate burst errors according to [27, 37, 22]. Furthermore, the QETD can also decode degenerate errors belong to the coset of $C^{\bot}$. In addition, QETD in Algorithm 3 is also available for Hermitian-ype QCCs. In Table IV, we list the numerical results of several QCCs for correcting quantum burst errors. Denote $\mathcal{N}_{D}$, $\mathcal{N}_{0}$, and $\mathcal{N}$ by the numbers of all correctable burst errors, all correctable nondegenerate burst errors and total errors of length $n\leq(n-\mathcal{K})/2$, respectively. We exhaustively traverse all the burst errors of length $\ell\leq(n-\mathcal{K})/2$ and count the numbers of burst errors that are successfully decoded by QETD. It is shown in Table IV that QETD can correct more degenerate burst errors than nondegenerate ones. As the code length grows, the ratio $\mathcal{N}_{D}/\mathcal{N}_{0}$ becomes extremely larger and thus QETD can correct much more degenerate burst errors than nondegenerate ones. As an example, nondegenerate burst errors that can be decoded by the $Q=[[29,1]]$ QCC are only $4.4\%$ of the total burst errors of length $\ell\leq 12$. However, $Q$ can decode degenerate burst errors up to $26.9\%$ of the total burst errors of length $\ell\leq 12$.
@@
$[[5,1]]$ $15$ $15$ $51$ $\% 29.429.4\%$ 11 $g=(\textbf{1}^{2}\textbf{2}^{1}\textbf{1}^{0})$
$[[7,1]]$ $72$ $57$ $255$ $8.2\% 22$$2.3\%$ $1.26$ $g=(\textbf{1}^{3}\textbf{1}^{1}\textbf{1}^{0})$
@@
$[[17,1]]$ $1.45401\rm{E}{5}$ $4.1064\rm{E}{4}$ $5.07903\rm{E}{5}$ $28.6\%$ $8\%$ $3.54$ $g=(\textbf{1}^{8}\textbf{3}^{7}\textbf{3}^{5}\textbf{3}^{4}\textbf{3}^{3}\textbf{3}^{1}\textbf{1}^{0})$
$[[23,1]]$ $1.1514471\rm{E}{7}$ $2.395308\rm{E}{6}$ $4.1943039\rm{E}{7}$ $27.4\%$ $5.7\%$ $4.81$ $g=(\textbf{1}^{11}\textbf{1}^{9}\textbf{1}^{7}\textbf{1}^{6}\textbf{1}^{5}\textbf{1}^{1}\textbf{1}^{0})$
$[[25,1]]$ $4.9269693\rm{E}{7}$ $9.363588\rm{E}{6}$ $1.80355071\rm{E}{8}$ $27.3\%$ $5.1\%$ $5.26$ $g=(\textbf{1}^{12}\textbf{2}^{11}\textbf{1}^{10}\textbf{2}^{7}\textbf{3}^{6}\textbf{2}^{5}\textbf{1}^{2}\textbf{2}^{1}\textbf{1}^{0})$
$[[29,1]]$ $8.86214133\rm{E}{8}$ $1.44826293\rm{E}{8}$ $3.288334248\rm{E}{9}$ $26.9\%$ $4.4\%$ $6.12$ $g=(\textbf{1}^{14}\textbf{2}^{13}\textbf{2}^{11}\textbf{3}^{10}\textbf{1}^{9}\textbf{3}^{8}\textbf{2}^{7}\textbf{3}^{6}\textbf{1}^{5}\textbf{3}^{4}\textbf{2}^{3}\textbf{2}^{1}\textbf{1}^{0})$